Q6. The lifetimes of two components in a machine have the following joint pdf: f(x, y).00-x y) for 0<50-y < 50 and zero elsewhere a. What is the probability that both components are functioning 20 months from now. b. What is the probability the component with life time X would fail 3 months before the other one? c. Compute the covariance of X, Y d. Compute the expected life of the machine e. What is probability that the two components fail within three months apart?
Answer #1

Solution : ( a )

P(X > 20, Y > 20)

-30 ,50- 125000(50 -y)dyda

From (1);

6 125000 ( - _ y) dy 20

Compute the indefinite integral 6 125000-y) dy

6 000 (50 -y) dy

Take the constant out : a * f (x) dx f (x) dr a *

6 125000 * 50-х

(50 - r- y)dy 62500

. J _jrr dtjg Apply the Sunn Rue (r) ±g(z) dr ) (r)dr

=rac{3}{62500}left(int :50dy-int :xdy-int :ydy ight)

=rac{3}{62500}left(50y-xy-rac{y^2}{2} ight)

Add a constant to the solution

62500 у-гу-

mathrm{Compute:the:boundaries}:quad int _{20}^{50-x}rac{6}{125000}left(50-x-y ight)dy

50-х 6 125000 (50-x-y) dy = y-linr-(02500 J20 6250050- , у-ху

=rac{3}{62500}left(50left(50-x ight)-xleft(50-x ight)-rac{left(50-x ight)^2}{2} ight)-rac{3}{62500}left(50*20-x*20-rac{20^2}{2} ight)

=rac{3left(50left(50-x ight)-xleft(50-x ight)-rac{left(50-x ight)^2}{2} ight)}{62500}-rac{3left(50*20-x*20-rac{(20)^2}{2} ight)}{62500}

2-100r +2500 62500 3 (1000-204) 62500

(2100 2500 3 (-20 1000- 200) 2 62500 62500

=rac{3left(x^2-100x+2500 ight)}{125000}-rac{3left(-20x+800 ight)}{62500}

=rac{3left(x^2-100x+2500 ight)}{125000}-rac{3left(-x+40 ight)}{3125}

3 (12-100x + 2500) 125000 100x + 250 --( 312540)

mathrm{Simplify}

=rac{27}{1250}+rac{3x^2-180x}{125000}

(50 -x -y)dy250 25000 312-180 :--→ (2) So, 125000 (50-x-y) dy = + 一→ (2) 20

mathrm{Substitute:(2):in:(1);}

=int _{20}^{30}left(rac{27}{1250}+rac{3x^2-180x}{125000} ight)dx

int _{20}^{30}left ( rac{27}{1250}+rac{3x^2-180x}{125000} ight )dx

27 3a2-180r 27 3r2180.r Compute the indefinite integral 1250 125000

27 3r2180.r 1250125000

. J _jrr dtjg Apply the Sunn Rue (r) ±g(z) dr ) (r)dr

=int rac{27}{1250}dx+int rac{3x^2-180x}{125000}dx

27 1250125000*/(87* - 180r) da (322 -180x) dr 1250 125000

27 2180dr 3r dr 1250 125000

=rac{27}{1250}x+rac{1}{125000}left(x^3-90x^2 ight)

Add a constant to the solution

=rac{27}{1250}x+rac{1}{125000}left(x^3-90x^2 ight)+C

mathrm{Compute:the:boundaries}:quad int _{20}^{30}left(rac{27}{1250}+rac{3x^2-180x}{125000} ight)dx

int _{20}^{30}left(rac{27}{1250}+rac{3x^2-180x}{125000} ight)dx=lim _{x o :30-}left(rac{27}{1250}x+rac{1}{125000}left(x^3-90x^2 ight) ight)-lim _{x o :20+}left(rac{27}{1250}x+rac{1}{125000}left(x^3-90x^2 ight) ight)

= (1270. 30+ 3030(30)2) 2) )-( 1250 * 20 + 125000 ((20)3-90 * (20)2) ) (30)- 9 1250 * 30 + 125000 (

=left ( rac{81}{125}-rac{54}{125} ight )-left ( rac{54}{125}-rac{28}{125} ight )

81-54 125 54-28 125

27 26 125 125

mathrm{Simplify}

125

0,008

ΓΓ 12600(50-1-V) dydr-125 Hence, (Decimal 0.008)